OpenStudy (rea__):

-7/3n = -42 n= ? EXPLANATION WOULD BE NICE

OpenStudy (anonymous):

Let's solve your equation step-by-step. −73n=−42 Step 1: Multiply both sides by 3/(-7). (3−7)*(−73n)=(3−7)*(−42) n=18 Answer: n=18

OpenStudy (anonymous):

the ones lke 73 is supposed to be 7/3

OpenStudy (anonymous):

@rea__

OpenStudy (anonymous):

Any ways Medal!! please :)

OpenStudy (anonymous):

TY :) for medal

OpenStudy (rea__):

@sleepyyy you were right thank u

OpenStudy (anonymous):

Np Lol any time you need help tell me @rea__ :)

OpenStudy (rea__):

@sleepyyy i do need help with a few